Need help solving for the "addition to retained earnings" question. Please show work and answer. Thank you
Consider the following income statement for the Heir Jordan Corporation: HEIR JORDAN CORPORATION Income Statement Sales Costs $49,000 40.300 Taxable income Taxes (22%) $ 8,700 1,914 Net income $ 6,786 Dividends Addition to retained4,386 $2.400 earnings A 20 percent growth rate in sales is projected. Prepare a pro forma income statement assuming costs vary with sales and the dividend payout ratio is constant (Input all answers as positive values. Do not round Intermedlate celculations.) HEIR JORDAN CORPORATION Pro Forma Income Statement 58.800 48.380 10.440 2 297 8.143 Saes Costs Net income
What Is the projected addition to retalned earnings? (Do not round Intermedlate to
Answer #1

please find below answer. .

Dividend payout ratio = divided/net income
=2400/6786
35.37%
Computation of net income statement
Sales 58800
Cost 48360
Taxable income 10440
Taxes 2297
Net income 8143
Computation of addition to retained earning
a Net income = 8143
b Dividend payout ratio = 35.37%
c=a*b Dividend paid = 2880
d=a-c addition to retained earning = 5263